# Gravity as a topological gauge theory

### v1

@@ Abstract
We describe a topological gauge theory with underlying Cartan geometry $G/H$ and principal $H$-bundle. The topological action is constructed as a linear combination of the Euler and Pontrjagin numbers coming from entries of the curvature $\bar{\Omega}$ of a Cartan connection. As an example we show that for respectively $G=\left\{\begin{array}[]{ll}SO(4,1)\text{ for }\Lambda_{0}>0\\ SO(3,2)\text{ for }\Lambda_{0}<0\end{array}\right.$ and $H=SO(3,1)$ a very particular linear combination yields the Holst + Euler and Pontrjagin of the curvature $R$ of the spin connection + Nieh-Yan + bare cosmological constant $\Lambda_{0}$ terms. Additionally, in this construction, the coupling constants of these different terms are inherently linked together. The other treated example is the one of the Möbius group $G=SO(4,2)/\{\pm I\}$ and $H=CO(4,2)\ltimes\mathbb{R}^{3,1}$. By considering a constraint on the "pair of frames" $(\alpha,\beta)$ we retrieve an action comprising all terms compatible with diffeomorphism invariance and local Lorentz invariance described in the previous example + kinetic term for a scalar field $+$ an interaction of that scalar field with the Torsion $T$, the scalar field being related to dilation. Finally, in both examples, we study the equations of motion associated to the total (Gauge + Matter) action. It is especially shown that one of the solutions in the Möbius case consists of Einstein’s equations modified by the Holst and Nieh-Yan terms with an additional source term for curvature depending on specific variations of spin density of matter.
@@ 1 Introduction
There exist several approaches to describe gravity as a gauge theory. Two of them are MacDowell-Mansouri gravity and BF theory [17]. They can be mathematically described in terms of Cartan geometries $G/H$ of manifolds $\mathscr{M}$ where $G,H\subset G$ and $\mathfrak{g},\mathfrak{h}$ are respectively Lie groups and their associated Lie algebras. The associated $\mathfrak{g}$-valued Cartan connection $\varpi$ naturally splits into two parts $\varpi=\varpi_{\mathfrak{h}}\oplus\varpi_{\mathfrak{p}}$ when the geometry is reductive, one with values in $\mathfrak{h}$ and one in $\mathfrak{p}=\mathfrak{g}/\mathfrak{h}$. In these formalisms the $\mathfrak{h}$-valued part $\varpi_{\mathfrak{h}}$ corresponds to the spin connection $A$ while the $\mathfrak{p}$-valued part $\varpi_{\mathfrak{p}}=\sqrt{\dfrac{\Lambda}{3\epsilon}}\,\beta$ is proportional to the vielbein 1-form $\beta$ (tetrad if $dim(\mathscr{M})=4$). For a resulting Lorentzian [^1] Riemaniann or Galilean spacetimes can also be described by this formalism. For more details see [17, p10]. The kind of topological gauge theories developed in this article are also compatible with these choices of Lie groups. space-time the choice of Lie groups is given by $H=SO(3,1)$ while $G$ is chosen according to the value and sign of the cosmological constant $\Lambda$ as in [17, p10]:
$\left\{\begin{array}[]{lcll}G=SO(4,1)&\text{ and }&\epsilon=+1,&\text{ for }\Lambda>0\\ G=SO(3,2)&\text{ and }&\epsilon=-1,&\text{ for }\Lambda<0\end{array}\right.$ (3)
@@
Let $\mathscr{M}$ be a manifold, $G$ a Lie group and $H$ a closed subgroup of $G$ such that $dim(G/H)=dim(\mathscr{M})=m$ and $\mathscr{P}$ a principal $H$-bundle over $\mathscr{M}$. Let $\mathfrak{X}\in\mathfrak{h}$, we denote $\mathfrak{X}^{v}$ the associated vertical vector field on $\mathscr{P}$. A Cartan connection on $\mathscr{P}$ is equivalent to a $1$-form $\varpi$ on $\mathscr{P}$ with values in $\mathfrak{g}$ the Lie algebra of $G$ satisfying the conditions [9, p128]:
$(a)~{}\varpi(\mathfrak{X}^{v})=\mathfrak{X},~{}\text{ for any }\mathfrak{X}\in\mathfrak{h}.$ (4)
$(b)~{}(R_{h})^{*}\varpi=Ad(h^{-1})\varpi~{},\forall h\in H,\text{where }R\text{ is the right action of }H\text{ on }\mathscr{P}$ (5)
@@
$(c)~{}\varpi(\mathfrak{X})\neq 0\text{ for every nonzero vector field }\mathfrak{X}\in\Gamma(T\mathscr{P}).$ (6)
Keywords : one, two, three, four
Let $\bar{\Omega}=d\varpi+\varpi\wedge\varpi=\bar{\Omega}_{\mathfrak{h}}\oplus\bar{\Omega}_{\mathfrak{p}}$ be the curvature of the Cartan connection $\varpi=\varpi_{\mathfrak{h}}\oplus\varpi_{\mathfrak{p}}$ and $\star$ an internal Hodge as used in [12] such that $\star\kappa^{ab}=\tfrac{1}{2}\varepsilon^{ab}_{~{}~{}kl}\kappa^{kl}$. For $m=4$, the MacDowell-Mansouri action is given by:
@@
This action is $Ad(H)$-invariant, i.e. Lorentz invariant. On the other hand, the Euler characteristic class $E(\mathscr{M})$ of a manifold $\mathscr{M}$ is null if $\dim(\mathscr{M})=m$ is odd, otherwise it can be linked to the $\tfrac{m}{2}$-th Pontrjagin class $P_{m/2}$ (more details can be found in [13]).
@@
$E(\mathscr{M})=\operatorname{Pf}(\dfrac{\Omega_{\mathfrak{so}}}{2\pi})$
@@
In particular, for the case of the Euler class for $m=4$, the Euler number is thus:
@@
Up to an overall factor, one can recognise the MacDowell-Mansouri action (7) is recast as the Euler number associated to the $\mathfrak{h}=\mathfrak{so}(3,1)$-valued part $\bar{\Omega}_{\mathfrak{h}}$ of the curvature and which is $SO(3,1$-invariant. This provides a very clear mathematical interpretation to the MacDowell-Mansouri action:
@@
Another characteristic class which will be of interest in the remainder of the article is the Pontrjagin characteristic class. As described in [13], one has the
@@
Let $L$ be a $k$-dimensional real vector bundle over an $m$-dimensional manifold $\mathscr{M}$. $\Omega$ is the curvature associated to a connection $\omega$ on the principal $H$-bundle over $\mathscr{M}$. Introducing orthonormal frames allows to reduce the structure group to $O(k)$ and one has $\mathfrak{o}(k)=\mathfrak{so}(k)$ at the level of Lie algebras. The $\mathfrak{so}(k)$-valued curvature is then $\Omega_{\mathfrak{so}}$. The corresponding total Pontrjagin class is then defined as: $P(\Omega_{\mathfrak{so}})=\det\Bigl{(}\mathds{1}+\dfrac{\Omega_{\mathfrak{so}}}{2\pi}\Big{)}$.
@@
where $P_{1}$ is the 1st Pontrjagin class. As stated in [13, Chern-Weil theorem], a nice feature of characteristic classes is that up to an exact term, they do not depend on the choice of connection.
Section 2 will introduce and describe the topological gauge theories associated to the Lie groups $G=SO(4,1),SO(3,2)$ considered as "mutations" of the groups $G=ISO(3,1)$. All these three groups contain $H=SO(3,1)$ which will be the structure group of the principal bundle $\mathscr{P}$. Then a topological action as a linear combination of the Euler and Pontrjagin numbers will be considered for these Lie groups $G$. Results of Section 2 are easily applicable to other spacetimes whether Riemaniann or Galilean or entirely different Lie groups. In particular, Section 3 will be devoted to the construction of a similar topological action for the case where $G=SO(4,2)/\{\pm I\}$ (the Möbius group). In the following, latin and greek letters will respectively be used for algebraic and space-time indices.
@@ 2.1 Lorentzian (dS and AdS) spacetimes as mutations of an
Let $G$, $H\subset G$ be Lie groups with corresponding Lie algebras $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{p}$ such that $G/H$ is reductive, that is that the splitting $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{p}$ is $Ad(H)$-invariant. One such reductive model geometry is $G/H=ISO(3,1)/SO(3,1)$, with algebra $\mathfrak{g}=\overbrace{\mathfrak{so}(3,1)}^{\mathfrak{h}}\oplus\overbrace{\mathfrak{\mathbb{R}}^{3,1}}^{\mathfrak{p}}$ where $G=ISO(3,1)=SO(3,1)\ltimes\mathbb{R}^{3,1}$ with $H=SO(3,1)$ defined by the relation $h^{\text{T}}\eta h=\eta~{},\forall h\in H$ where $\eta$ is the Minkowski metric:
@@
A group element $g\in G$ is given by :
$g=\begin{pmatrix}\Lambda&\dfrac{1}{\ell}v\\ 0&1\end{pmatrix}\in G\text{ where }\Lambda\in SO(3,1)\text{ such that }\Lambda^{\text{T}}\eta\Lambda=\eta\text{ and }v\in\mathbb{R}^{(3,1)}.$
@@
with $\{J_{ab}\}_{a,b=0,...,3}$ generators of $\mathfrak{h}=\mathfrak{so}(3,1)$ and $\{e_{a}\}_{a=0,...,3}$ a basis of $\mathfrak{p}=\mathbb{R}^{3,1}$. In [16, p218], Sharpe introduces the notion of mutation of a Cartan geometry $G/H$ according to the
@@
with $k$ a 0-form scalar field and $\varphi_{\bar{\mathfrak{p}}}=\varphi_{\mathfrak{p}}^{\text{T}}\eta$. One can observe that $(\mathfrak{g}^{\prime},\mathfrak{h})$ reduces to dS and AdS geometries respectively for the two particular cases, see e.g. [17]:
@@
Throughout the paper $k$ will be considered as a real number, the case of a 00-form scalar field $k$ will be studied in a forthcoming paper. On the other hand, the canonical decomposition of a symmetric algebra $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$ (see [11, Prop.2.1, Chapter XI], and [10, chapter 11] for more examples) corresponds to the commutators:
@@
Therefore, one can see that the mutated Lie algebra $\mathfrak{g}^{\prime}=\mathfrak{h}\oplus\mathfrak{m}$ of (5) corresponds to a symmetric algebra for $k\neq 0$, as a particular case of a reductive geometry with $\mathfrak{m}=\mathfrak{g}^{\prime}/\mathfrak{h}\simeq\mathfrak{p}=\mathfrak{g}/\mathfrak{h}$. Moreover, the generators of $\mathfrak{h}=\mathfrak{so}(3,1)$, and elements of the bases of $\mathfrak{p}=\mathfrak{\mathbb{R}}^{3,1}$, $\mathfrak{p}^{*}=\mathfrak{\mathbb{R}}^{*3,1}$ and ${\bar{\mathfrak{p}}}=\bar{\mathbb{R}}^{3,1}$ obey the relations (the isomorphism $\mathrm{End}(\mathbb{R}^{n})\simeq\mathbb{R}^{n}\otimes{\mathbb{R}^{n}}^{*}$ is used):
@@
with $C^{fe}_{ab,cd}=-C^{ef}_{ab,cd}$ which is easily computed. An element $\varphi=\varphi_{\mathfrak{h}}\oplus\varphi_{\mathfrak{m}}\in\mathfrak{g}^{\prime}$ admits the following matrix representation:
@@
with $\varphi^{a}_{\mathfrak{m}}=\varphi^{a}_{\mathfrak{p}}$ and $M_{a}=\begin{pmatrix}0&e_{a}\\ k\bar{e}_{a}&0\end{pmatrix}$ the generators of $\mathfrak{m}$. Let $\varphi,\varphi^{\prime}\in\mathfrak{g}^{\prime}$, their Lie bracket reads:

### v3

@@ Abstract
We describe a topological gauge theory with underlying Cartan geometry $G/H$ and principal $H$-bundle. The topological action is constructed as a linear combination of the Euler and Pontrjagin numbers coming from entries of the curvature $\bar{\Omega}$ of a Cartan connection.
As an example we show that for $G=SO(4,1)$ (respectively $SO(3,2)$) for $\Lambda_{0}>0$ (respectively $\Lambda_{0}<0$), and $H=SO(3,1)$, a very particular linear combination yields the Holst + Euler and Pontrjagin of the curvature $R$ of the spin connection + Nieh-Yan + bare cosmological constant $\Lambda_{0}$ terms. Additionally, in this construction, the coupling constants of these different terms are inherently linked together.
The other example treated is the one of the Möbius group $G=SO(4,2)/\{\pm I\}$ and $H=CO(4,2)\ltimes\mathbb{R}^{*3,1}$. By considering a constraint on the "pair of frames" $(\alpha,\beta)$ we retrieve an action comprising all terms compatible with diffeomorphism invariance and local Lorentz invariance described in the previous example + kinetic term for a scalar field $+$ an interaction of that scalar field with the Torsion $T$, the scalar field being related to dilations. Finally, in both examples, we study the equations of motion associated to the total (Gauge + Matter) action. It is especially shown that one of the solutions in the Möbius case consists of Einstein’s equations modified by the Holst and Nieh-Yan terms with an additional source term for curvature depending on specific variations of spin density of matter and on torsion.
@@ 1 Introduction
There exist several approaches to describe gravity as a gauge theory. Two of them are MacDowell-Mansouri gravity and BF theory, see e.g. [1]. They can be mathematically described in terms of Cartan geometries $G/H$ of manifolds $\mathscr{M}$ where $G,H\subset G$ and $\mathfrak{g},\mathfrak{h}$ are respectively Lie groups and their corresponding Lie algebras. When the geometry is reductive, the associated $\mathfrak{g}$-valued Cartan connection $\varpi$ naturally splits into two parts $\varpi=\varpi_{\mathfrak{h}}\oplus\varpi_{\mathfrak{p}}$, one with values in $\mathfrak{h}$ and the other one in $\mathfrak{p}=\mathfrak{g}/\mathfrak{h}$. In these formalisms the $\mathfrak{h}$-valued part $\varpi_{\mathfrak{h}}$ corresponds to the spin connection $A$ while the $\mathfrak{p}$-valued part $\varpi_{\mathfrak{p}}=\sqrt{\dfrac{\Lambda}{3\epsilon}}\,\beta$ is proportional to the vielbein 1-form $\beta$ (tetrad if $dim(\mathscr{M})=4$).
For a resulting Lorentzian [^3] Riemaniann or Galilean spacetimes can also be described by this formalism. For more details see [1, p.10]. The kind of topological gauge theories developed in this article are also compatible with these choices of Lie groups. space-time the structure group is given by $H=SO(3,1)$ while the Lie group $G$ is chosen according to the value and sign of the cosmological constant $\Lambda$ as in [1, p.10]:
$\left\{\begin{array}[]{lcll}G=SO(4,1)&\text{ and }&\epsilon=+1,&\text{ for }\Lambda>0\\ G=SO(3,2)&\text{ and }&\epsilon=-1,&\text{ for }\Lambda<0.\end{array}\right.$ (3)
Let us recall the
@@
Let $\mathscr{M}$ be a manifold, $G$ a Lie group and $H$ a closed subgroup of $G$ such that $dim(G/H)=dim(\mathscr{M})=m$ and $\mathscr{P}$ a principal $H$-bundle over $\mathscr{M}$. Let $\mathfrak{X}\in\mathfrak{h}$, we denote $\mathfrak{X}^{v}$ the associated vertical vector field on $\mathscr{P}$. A Cartan connection on $\mathscr{P}$ is equivalent to a $1$-form $\varpi$ on $\mathscr{P}$ with values in $\mathfrak{g}$ the Lie algebra of $G$ satisfying the conditions [2, p128]:
$(a)~{}\varpi(\mathfrak{X}^{v})=\mathfrak{X},~{}\text{ for any }\mathfrak{X}\in\mathfrak{h}.$
$(b)~{}(R_{h})^{*}\varpi=Ad(h^{-1})\varpi~{},\forall h\in H,\text{where }R\text{ is the right action of }H\text{ on }\mathscr{P}$
@@
$(c)~{}\varpi(\mathfrak{X})\neq 0\text{ for every nonzero vector field }\mathfrak{X}\in\text{Vect}(\mathscr{P}).$
Let $\bar{\Omega}=d\varpi+\varpi\wedge\varpi=\bar{\Omega}_{\mathfrak{h}}\oplus\bar{\Omega}_{\mathfrak{p}}$ be the curvature of the Cartan connection $\varpi=\varpi_{\mathfrak{h}}\oplus\varpi_{\mathfrak{p}}$ and $\star$ an internal Hodge as used in [3] such that $\star\kappa^{ab}=\tfrac{1}{2}\varepsilon^{ab}_{~{}~{}kl}\kappa^{kl}$.
For $m=4$, the MacDowell-Mansouri action is given by:
@@
This action is $Ad(H)$-invariant, i.e. Lorentz invariant.
On the other hand, the Euler characteristic class $E(\mathscr{M})$ of a manifold $\mathscr{M}$ is null if $\dim(\mathscr{M})=m$ is odd, otherwise it can be linked to the $\tfrac{m}{2}$-th Pontrjagin class $P_{m/2}$ (more details can be found in [4]).
@@
$E(\mathscr{M})=\operatorname{Pf}\left(\dfrac{\Omega_{\mathfrak{so}}}{2\pi}\right)$
@@
In particular, for $m=4$, the Euler number is thus:
@@
Up to an overall factor, one can recognise the MacDowell-Mansouri action (4) which is recast as the Euler number associated to the $\mathfrak{h}=\mathfrak{so}(3,1)$-valued part $\bar{\Omega}_{\mathfrak{h}}$ of the curvature and which is $SO(3,1)$-invariant. This provides a very clear mathematical interpretation to the MacDowell-Mansouri action:
@@
Another characteristic class which will be of interest in the remainder of the article is the Pontrjagin characteristic class. As described in [4], one has:
@@
Let $L$ be a $k$-dimensional real vector bundle over an $m$-dimensional manifold $\mathscr{M}$. $\Omega$ is the curvature associated to a connection $\omega$ on the principal $H$-bundle over $\mathscr{M}$. Introducing orthonormal frames allows to reduce the structure group to $O(k)$ and one has $\mathfrak{o}(k)=\mathfrak{so}(k)$ at the level of Lie algebras. The $\mathfrak{so}(k)$-valued curvature is written $\Omega_{\mathfrak{so}}$. The corresponding total Pontrjagin class is then defined as: $P(\Omega_{\mathfrak{so}})=\det\Bigl{(}\mathds{1}+\dfrac{\Omega_{\mathfrak{so}}}{2\pi}\Big{)}$.
@@
where $P_{1}$ is the $1$st Pontrjagin class.
As stated in [4, Chern-Weil theorem], a nice feature of characteristic classes is that up to an exact term, they do not depend on the choice of connection.
Section 2 will introduce and describe the topological gauge theories associated to the Lie groups $G=SO(4,1),SO(3,2)$ considered as "mutations" (according to [5]) of the group $G=ISO(3,1)$. All these three groups contain $H=SO(3,1)$ which will be the structure group of the principal bundle $\mathscr{P}$. Then, a topological action as a linear combination of the Euler and Pontrjagin numbers will be considered for these Lie groups $G$ and $H$. Results of Section 2 are easily applicable to other spacetimes whether Riemaniann or Galilean or entirely different Lie groups.
In particular, Section 3 will be devoted to the construction of a similar topological action for the case where $G=SO(4,2)/\{\pm I\}$ (the Möbius group) and $H=CO(4,2)\ltimes\mathbb{R}^{*3,1}$, with $CO(p,q)$ the conformal group of signature $(p,q)=(4,2)$ here.
In the following, latin and greek letters will respectively be used for algebraic and space-time indices.
@@ 2.1 Lorentzian (dS and AdS) spacetimes as mutations of an
Let $G$, $H\subset G$ be Lie groups with corresponding Lie algebras $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{p}$ such that $G/H$ is reductive, that is that the splitting $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{p}$ is $Ad(H)$-invariant.
One such reductive model geometry is $G/H=ISO(3,1)/SO(3,1)$, with algebra $\mathfrak{g}=\overbrace{\mathfrak{so}(3,1)}^{\mathfrak{h}}\oplus\overbrace{\mathfrak{\mathbb{R}}^{3,1}}^{\mathfrak{p}}$ where $G=ISO(3,1)=SO(3,1)\ltimes\mathbb{R}^{3,1}$ and $H=SO(3,1)$ is defined by the relation
$\Lambda^{\text{T}}\eta\Lambda=\eta,\forall\Lambda\in SO(3,1)$ with $\eta$ the Minkowski metric:
@@
A group element $g\in G=ISO(3,1)$ is parametrized as ${\displaystyle g=\begin{pmatrix}\Lambda&\dfrac{1}{\ell}v\\ 0&1\end{pmatrix}}$ where $v\in\mathbb{R}^{3,1}$.
@@
with $\{J_{ab}\}_{a,b=0,...,3}$ the generators of $\mathfrak{h}=\mathfrak{so}(3,1)$ and $\{e_{a}\}_{a=0,...,3}$ a basis of $\mathfrak{p}=\mathbb{R}^{3,1}$.
In [5, p218], Sharpe introduces the notion of mutation of a Cartan geometry $G/H$ according to the
@@
with $\varphi_{\bar{\mathfrak{p}}}=\varphi_{\mathfrak{p}}^{\text{T}}\eta$ and where $k$ may a priori be considered as 00-form scalar field.
One can observe that $(\mathfrak{g}^{\prime},\mathfrak{h})$ reduces to dS and AdS geometries respectively for the two particular cases, see e.g. [1]:
@@
Throughout the paper $k$ will be considered as a real number, the case of a 00-form scalar field $k$ will be studied in a forthcoming paper.
On the other hand, the canonical decomposition of a symmetric algebra $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$ (see [6, Prop.2.1, Chapter XI] for more examples) corresponds to the commutators:
@@
Therefore, one can see that the mutated Lie algebra $\mathfrak{g}^{\prime}=\mathfrak{h}\oplus\mathfrak{m}$ of (1) corresponds to a symmetric algebra for $k\neq 0$, as a particular case of a reductive geometry with $\mathfrak{m}=\mathfrak{g}^{\prime}/\mathfrak{h}\simeq\mathfrak{p}=\mathfrak{g}/\mathfrak{h}$. Moreover, the generators of $\mathfrak{h}=\mathfrak{so}(3,1)$, and elements of the basis of $\mathfrak{p}=\mathfrak{\mathbb{R}}^{3,1}$, $\mathfrak{p}^{*}=\mathfrak{\mathbb{R}}^{3,1*}$ and ${\bar{\mathfrak{p}}}=\bar{\mathbb{R}}^{3,1}$ obey the relations (the isomorphism $\mathrm{End}(\mathbb{R}^{n})\simeq\mathbb{R}^{n}\otimes{\mathbb{R}^{n*}}$ is used):
@@
with $C^{fe}_{ab,cd}=-C^{ef}_{ab,cd}$ which is easily computed.
An element $\varphi=\varphi_{\mathfrak{h}}\oplus\varphi_{\mathfrak{m}}\in\mathfrak{g}^{\prime}=\mathfrak{h}\oplus\mathfrak{m}$ admits the following matrix representation:
@@
where $J_{ab}$ is canonically identified with the matrix representation of $\mathfrak{h}\subset\mathfrak{g}$, $\varphi^{a}_{\mathfrak{m}}=\varphi^{a}_{\mathfrak{p}}$ and $M_{a}=\begin{pmatrix}0&e_{a}\\ k\bar{e}_{a}&0\end{pmatrix}$ the generators of $\mathfrak{m}$.
Let $\varphi,\varphi^{\prime}\in\mathfrak{g}^{\prime}=\mathfrak{h}\oplus\mathfrak{m}$, their Lie bracket is given by the commutator and reads: